# A Survey of Personalized Large Language Models: Progress and Future Directions

### v1

@@ 3.3 Soft-Fused Prompting
Soft prompting differs from profile-augmented prompting by compressing personalized data into soft embeddings, rather than summarizing it into discrete tokens. These embeddings are generated by a user feature encoder .
@@
Soft prompting, used as an input prefix, focuses on the embedding level by concatenating the query embedding with the soft embedding, and is commonly applied in recommendation tasks. UEM (Doddapaneni et al., 2024) is a user embedding module (transformer network) that generates a soft prompt conditioned on the user’s personalized data. PERSOMA (Hebert et al., 2024) enhances UEM by employing resampling, selectively choosing a subset of user interactions based on relevance and importance. REGEN (Sayana et al., 2024) combines item embeddings from user-item interactions via collaborative filtering and item descriptions using a soft prompt adapter to generate contextually personalized responses. PeaPOD (Ramos et al., 2024) personalizes soft prompts by distilling user preferences into a limited set of learnable, dynamically weighted prompts. Unlike previously mentioned methods, which focus on directly embedding user interactions or resampling relevant data, PeaPOD adapts to user interests by weighting a shared set of prompts.
Cross-Attention
@@
Cross-attention enables the model to process and integrate multiple input sources by allowing it to attend to personalized data and the query. User-LLM (Ning et al., 2024) uses an autoregressive user encoder to convert historical interactions into embeddings through self-supervised learning, which are then integrated via cross-attention. The system employs joint training to optimize both the retriever and generator for better performance. RECAP (Liu et al., 2023) utilizes a hierarchical transformer retriever designed for dialogue domains to fetch personalized information. This information is integrated into response generation via a context-aware prefix encoder, improving the model’s ability to generate personalized, contextually relevant responses.
Output Logits
@@
GSMN (Wu et al., 2021) retrieves relevant information from personalized data, encodes it into soft embeddings, and uses them in attention with the query vector. Afterward, the resulting embeddings are concatenated with the LLM-generated embeddings, modifying the final logits to produce more personalized and contextually relevant responses.
@@ 3.4 Discussions
The three prompting methods have distinct pros and cons: 1) Profile-augmented prompting improves efficiency by compressing historical data but risks information loss and reduced personalization. 2) Retrieval-augmented prompting offers rich, context-aware inputs and scales well for long-term memory but can suffer from computational limits and irrelevant data retrieval. 3) Soft prompting efficiently embeds user-specific info, capturing semantic nuances without redundancy, but is limited to black-box models and lacks explicit user preference analysis. Overall, prompting-based methods are efficient and adaptable, and enable dynamic personalization with minimal computational overhead. However, they lack deeper personalization analysis as they rely on predefined prompt structures to inject user-specific information and are limited in accessing global knowledge due to the narrow scope of prompts.
@@ 4 Personalized Adaptation
PLLMs require balancing fine-tuning’s deep adaptability with the efficiency of prompting. Therefore, specialized methods need to be specifically designed for PLLMs to address these challenges utilizing parameter-efficient fine-tuning methods (PEFT), such as LoRA (Hu et al., 2021; Yang et al., 2024), IA3 (Liu et al., 2022), etc. (See Figure 4).
@@ 4.1 One PEFT All Users
Single PEFT
PLoRA (Zhang et al., 2024c) and LM-P (Woźniak et al., 2024) utilize LoRA for PEFT of LLM, injecting personalized information via user embeddings and user IDs, respectively. PLoRA is further extended and supports online training and prediction for cold-start scenarios. UserIdentifier (Mireshghallah et al., 2021) uses a static, non-trainable user identifier to condition the model on user-specific information, avoiding the need for trainable user-specific parameters and reducing training costs. Review-LLM (Peng et al., 2024b) aggregates users’ historical behaviors and ratings into prompts to guide sentiment and leverages LoRA for efficient fine-tuning. However, these methods rely on a single architecture with fixed configurations (e.g., hidden size, insertion layers), making them unable to store and activate diverse information for personalization (Zhou et al., 2024). To solve this problem, MiLP (Zhang et al., 2024b) utilizes a Bayesian optimization strategy to automatically identify the optimal configuration for applying multiple LoRA modules, enabling efficient and flexible personalization.
Mixture of Experts (MoE)
Several methods use the LoRA module, but with a static configuration for all users. This lack of parameter personalization limits adaptability to user dynamics and preference shifts, potentially resulting in suboptimal performance (Cai et al., 2024). RecLoRA (Zhu et al., 2024) addresses this limitation by maintaining a set of parallel, independent LoRA weights and employing a soft routing method to aggregate meta-LoRA weights, enabling more personalized and adaptive results. Similarly, iLoRA (Kong et al., 2024) creates a diverse set of experts (LoRA) to capture specific aspects of user preferences and generates dynamic expert participation weights to adapt to user-specific behaviors.
@@ 4.2 One PEFT Per User
No Collaboration
There is no collaboration or coordination between adapters or during the learning process for each use in this category. UserAdapter (Zhong et al., 2021) personalizes models through prefix-tuning, fine-tuning a unique prefix vector for each user while keeping the underlying transformer model shared and frozen. PocketLLM (Peng et al., 2024a) utilizes a derivative-free optimization approach, based on MeZo (Malladi et al., 2023), to fine-tune LLMs on memory-constrained mobile devices. OPPU (Tan et al., 2024b) equips each user with a LoRA module.
Collaborative Efforts
The “one-PEFT-per-user” paradigm without collaboration is computationally and storage-intensive, particularly for large user bases. Additionally, individually owned PEFTs hinder community value, as personal models cannot easily share knowledge or benefit from collaborative improvements. PER-PCS (Tan et al., 2024a) enables efficient and collaborative PLLMs by sharing a small fraction of PEFT parameters across users. It first divides PEFT parameters into reusable pieces with routing gates and stores them in a shared pool. For each target user, pieces are autoregressively selected from other users, ensuring scalability, efficiency, and personalized adaptation without additional training.
@@ 4.3 Discussions
Fine-tuning methods enable deep personalization by modifying a large set of model parameters, and parameter-efficient fine-tuning methods (e.g., prefix vectors or adapters) reduce computational cost and memory requirements while maintaining high personalization levels. These methods improve task adaptation by tailoring models to specific user needs, enhancing performance in tasks like sentiment analysis and recommendations. They also offer flexibility, allowing user-specific adjustments while leveraging pre-trained knowledge. However, they still face the risk of overfitting, particularly with limited or noisy user data, which can impact generalization and performance for new or diverse users.
@@ 5 Personalized Alignment
5.1 Personalized Alignment Data Construction
High-quality data construction is critical for learning PLLMs, primarily involving self-generated data through interactions with the LLM. Wu et al. (2024c) constructs a dataset for aligning LLMs with individual preferences by initially creating a diverse pool of 3,310 user personas, which are expanded through iterative self-generation and filtering. This method is similar to PLUM (Magister et al., 2024) that both simulate dynamic interactions through multi-turn conversation trees, allowing LLMs to infer and adapt to user preferences. To enable LLMs to adapt to individual user preferences without re-training, Lee et al. (2024) utilizes diverse system messages as meta-instructions to guide the models’ behavior. To support this, the MULTIFACETED COLLECTION dataset is created, comprising 197k system messages that represent a wide range of user values. To facilitate real-time, privacy-preserving personalization on edge devices while addressing data privacy, limited storage, and minimal user disruption, Qin et al. (2024) introduces a self-supervised method that efficiently selects and synthesizes essential user data, improving model adaptation with minimal user interaction.
Research efforts are also increasingly concentrating on developing datasets that assess models’ comprehension of personalized preferences. Kirk et al. (2024) introduces PRISM Alignment Dataset that maps the sociodemographics and preferences of 1,500 participants from 75 countries to their feedback in live interactions with 21 LLMs, focusing on subjective and multicultural perspectives on controversial topics. PersonalLLM (Zollo et al., 2024) introduces a novel personalized testdb, which curates open-ended prompts and multiple high-quality responses to simulate diverse latent preferences among users. It generates simulated user bases with varied preferences from pre-trained reward models, addressing the challenge of data sparsity in personalization.
@@ 5.2 Personalized Alignment Optimization
Personalized preference alignment is usually modeled as a multi-objective reinforcement learning (MORL) problem, where personalized preference is determined as the user-specific combination of multi-preference dimensions. Based on this, a typical alignment paradigm involves using a personalized reward derived from multiple reward models to guide during the training phase of policy LLMs, aiming for personalization. MORLHF (Wu et al., 2023) separately trains reward models for each dimension and retrains the policy language models using proximal policy optimization, guided by a linear combination of these multiple reward models. This approach allows for the reuse of the standard RLHF pipeline. MODPO (Zhou et al., 2023) introduces a novel RL-free algorithm extending Direct Preference Optimization (DPO) for managing multiple alignment objectives. It integrates linear scalarization directly into the reward modeling process, allowing it to train language models via a simple margin-based cross-entropy loss as implicit collective rewards functions.
@@
Another strategy for MORL is to consider ad-hoc combinations of multiple trained policy LLMs during the decoding phase to achieve personalization. Personalized Soups (Jang et al., 2023) and Reward Soups (Rame et al., 2024) address the challenge of RL from personalized human feedback by first training multiple policy models with distinct preferences independently and then merging their parameters post-hoc during inference. Both methods allow for dynamic weighting of the networks based on user preferences, enhancing model alignment and reducing reward misspecification. Also, the personalized fusion of policy LLMs can be achieved not only through parameter merging but also through model ensembling. MOD (Shi et al., 2024) outputs the next token from a linear combination of all base models, allowing for precise control over different objectives by combining their predictions without the need for retraining. The method demonstrates significant effectiveness when compared to the parameter-merging baseline. PAD (Chen et al., 2024b) leverages a personalized reward modeling strategy to generate token-level personalized rewards, which are then used to guide the decoding process, allowing dynamic tailoring of the base model’s predictions to individual preferences. Figure 5 visualizes above typical approaches of MORL for personalized alignment.
There are some other emerging personalized alignment studies beyond the “multi-objective” paradigm. PPT (Lau et al., 2024) unlocks the potential of in-context learning for scalable and efficient personalization by generating two potential responses for each user prompt, asking the user to rank them, and incorporating this feedback into the model’s context to dyanmic adapt to individual preferences over time. VPL (Poddar et al., 2024) utilizes a variational inference framework to capture diverse human preferences through user-specific latent variables. Inferring user-specific latent distributions from a few preference annotations enables more accurate and personalized reward modeling with better data efficiency.

### v2

@@ 3.3 Soft-Fused Prompting
Soft prompting differs from profile-augmented prompting in that it compresses personalized data into soft embeddings rather than summarizing it into discrete tokens. These embeddings are generated by a user feature encoder .
@@
Input Prefix Soft prompting, used as an input prefix, focuses on the embedding level by concatenating the query embedding with the soft embedding, and is commonly applied in recommendation tasks. PPlug (Liu et al., 2024b) constructs a user-specific embedding for each individual by modeling their historical contexts using a lightweight plug-in user embedder module. This embedding is then attached to the task input. UEM (Doddapaneni et al., 2024) is a user embedding module (transformer network) that generates a soft prompt conditioned on the user’s personalized data. PERSOMA (Hebert et al., 2024) enhances UEM by employing resampling, selectively choosing a subset of user interactions based on relevance and importance. REGEN (Sayana et al., 2024) combines item embeddings from user-item interactions via collaborative filtering and item descriptions using a soft prompt adapter to generate contextually personalized responses. PeaPOD (Ramos et al., 2024) personalizes soft prompts by distilling user preferences into a limited set of learnable, dynamically weighted prompts. Unlike previously mentioned methods, which focus on directly embedding user interactions or resampling relevant data, PeaPOD adapts to user interests by weighting a shared set of prompts. ComMer (Zeldes et al., 2025) tunes an encoder to compress historical documents into embeddings, merges them through mean pooling, then feeds the result into a frozen base LLM, distinguishing it from methods focused on historical interaction data. DEP (Qiu et al., 2025b) further models inter-user differences in the latent space instead of merely considering the user’s own information.
@@
Cross-Attention Cross-attention enables the model to process and integrate multiple input sources by allowing it to attend to personalized data and the query. User-LLM (Ning et al., 2024) uses an autoregressive user encoder to convert historical interactions into embeddings through self-supervised learning, which are then integrated via cross-attention. The system employs joint training to optimize both the retriever and generator for better performance. RECAP (Liu et al., 2023) utilizes a hierarchical transformer retriever designed for dialogue domains to fetch personalized information. This information is integrated into response generation via a context-aware prefix encoder, improving the model’s ability to generate personalized, contextually relevant responses.
@@
Output Logits GSMN (Wu et al., 2021) retrieves relevant information from personalized data, encodes it into soft embeddings, and uses them in attention with the query vector. Afterward, the resulting embeddings are concatenated with the LLM-generated embeddings, modifying the final logits to produce more personalized and contextually relevant responses.
3.4 Contrastive Prompting
The key insight of contrastive prompting is to utilize two forward paths of LLM to generate contrast pairs with and without personalized information (See Figure 5 (d)). By comparing model states with and without personalization, it identifies factors influencing personalization, enabling dynamic adjustment of the model’s personalization level. Two mainstream model states are utilized: hidden states (representations) and logits (predictive distributions).
CoS (He et al., 2024) is a special case that assumes there is a brief user profile for each query; it amplifies its influence in LLM response generation by combining output probabilities (logits) with and without the profile, i.e., , adjusting personalization degree through hyperparameter without fine-tuning. In contrast, StyleVector (Zhang et al., 2025b) obtains encoded personalized information in LLM via hidden representations. Specifically, given a user and the personalized data (user historical content), , it use GPT-3.5-Turbo to generate a general response for comparison with the expected personalized response , i.e., , where extracts hidden representations in base model ’s middle layer. extracts the user style-based vector, and by adding this vector with a scaling factor , this method can steer the LLM toward personalized output to a controllable degree. CoSteer (Lv et al., 2025) performs decoding-time personalization through local delta steering. A local SLM produces logits with and without personal context, and their difference is used to steer the cloud LLM’s logits iteratively, preserving privacy while achieving controllable personalization without fine-tuning.
@@ 3.5 Discussions
The three prompting methods have distinct pros and cons (Table 2):
This approach centers on maintaining, selecting, and integrating external personalized data into LLMs to improve user behavior comprehension. Overall, prompting-based methods are efficient and adaptable, enabling dynamic personalization with minimal computational overhead. Especially for RAG, which reduces LLM hallucination by grounding responses in retrieved evidence. Therefore, prompting methods are particularly suitable for explicit queries that require answering specific factual information, as mentioned in Section 2.2, excelling when personalization involves retrieving concrete information from user data.
However, they lack deeper personalization analysis, as they rely on predefined prompt structures to inject user-specific information and are limited in accessing global knowledge due to the narrow scope of prompts, which usually fail in tasks with implicit generation queries compared with adaptation and alignment methods (Tan et al., 2024a).
@@ 4 Personalized Adaptation
| One PEFT for All Users (Shared Parameters, Section 4.1) | One PEFT Per User (Individual Parameters, Section 4.2) | |
| Training Objective | ||
| Inference Stage | ||
| Pros. | Parameter Efficiency Scalability | Strong Personalization User Isolation |
| Cons. | Limited Personalization Persnonalized Data Dependency | Storage Overhead Training Complexity |
PLLMs require balancing fine-tuning’s deep adaptability with the efficiency of prompting. Therefore, specialized methods need to be specifically designed for PLLMs to address these challenges utilizing parameter-efficient fine-tuning methods (PEFT), such as LoRA (Hu et al., 2021; Yang et al., 2024), prefix-tuning (Li and Liang, 2021), MeZo (Malladi et al., 2023), etc. From an architectural perspective, these methods can be categorized into two types: (1) one PEFT for all users, where all users share a single fine-tuning module that differentiates personalized information, and (2) one PEFT per user, where each user has their own unique fine-tuning block. Details are illustrated in Figure 6 and Table 3.
@@ 4.1 One PEFT All Users
Single PEFT PLoRA (Zhang et al., 2024d) and LM-P (Woźniak et al., 2024) utilize LoRA for PEFT of LLM, injecting personalized information via user embeddings and user IDs, respectively. PLoRA is further extended and supports online training and prediction for cold-start scenarios. UserIdentifier (Mireshghallah et al., 2021) uses a static, non-trainable user identifier to condition the model on user-specific information, avoiding the need for trainable user-specific parameters and reducing training costs. Review-LLM (Peng et al., 2024a) aggregates users’ historical behaviors and ratings into prompts to guide sentiment and leverages LoRA for efficient fine-tuning. However, these methods rely on a single architecture with fixed configurations (e.g., hidden size, insertion layers), making them unable to store and activate diverse information for personalization (Zhou et al., 2024). To solve this problem, MiLP (Zhang et al., 2024e) utilizes a Bayesian optimization strategy to automatically identify the optimal configuration for applying multiple LoRA modules, enabling efficient and flexible personalization.
Mixture of Experts (MoE) Several methods use the LoRA module, but with a static configuration for all users. This lack of parameter personalization limits adaptability to user dynamics and preference shifts, potentially resulting in suboptimal performance (Cai et al., 2024). RecLoRA (Zhu et al., 2024a) addresses this limitation by maintaining a set of parallel, independent LoRA weights and employing a soft routing method to aggregate meta-LoRA weights, enabling more personalized and adaptive results. Similarly, iLoRA (Kong et al., 2024) creates a diverse set of experts (LoRA) to capture specific aspects of user preferences and generates dynamic expert participation weights to adapt to user-specific behaviors.
@@ 4.2 One PEFT Per User
No Collaboration There is no collaboration or coordination between adapters or during the learning process for each use in this category. UserAdapter (Zhong et al., 2021) personalizes models through prefix-tuning, fine-tuning a unique prefix vector for each user while keeping the underlying transformer model shared and frozen. PocketLLM (Peng et al., 2024b) utilizes a derivative-free optimization approach, based on MeZo (Malladi et al., 2023), to fine-tune LLMs on memory-constrained mobile devices.
Collaborative Efforts The “one-PEFT-per-user" paradigm without collaboration is computationally and storage-intensive, particularly for large user bases. Additionally, individually owned PEFTs hinder community value, as personal models cannot easily share knowledge or benefit from collaborative improvements. OPPU (Tan et al., 2024a) frames the personalized LLMs training as a two-stage process, where a global LoRA module is learned to capture shared knowledge across users initially, followed by user-specific LoRA modules for individual adaptation. PerFit (Liu et al., 2025) further enhances this by replacing LoRA modules with representation fine-tuning modules, motivated by their finding that the personalization of LLM correlates with a low-rank collaborative shift and diverse personalized shifts following the collaborative one. CoPe (Bu et al., 2025) utilizes reward-guided decoding tailored for personalization, aiming to maximize the implicit reward signal for each user. PER-PCS (Tan et al., 2024b) enables efficient and collaborative PLLMs by sharing a small fraction of PEFT parameters across users. It first divides PEFT parameters into reusable pieces with routing gates and stores them in a shared pool. For each target user, pieces are autoregressively selected from other users, ensuring scalability, efficiency, and personalized adaptation without additional training. PROPER (Zhang et al., 2025c) introduces hierarchical user grouping with MoE-LoRA integration, enabling fine-grained personalization through staged adaptation from population-level, group-level, to individual-level models.
@@ 4.3 Discussions
Fine-tuning methods enable deep personalization by modifying a large set of model parameters, and parameter-efficient fine-tuning methods (e.g., prefix vectors or adapters) reduce computational cost and memory requirements while maintaining high personalization levels. These methods improve task adaptation by tailoring models to specific user needs, enhancing performance in tasks like sentiment analysis and recommendations. They also offer flexibility, allowing user-specific adjustments while leveraging pre-trained knowledge. However, they still face several challenges:
Addressing these issues calls for complementary strategies such as federated learning, synthetic data generation, and hybrid prompting–adaptation pipelines, in order to balance efficiency, personalization depth, and reliability.
@@ 5 Personalized Alignment
5.1 Data Construction
High-quality data construction is critical for learning PLLMs, primarily involving self-generated data through interactions with the LLM. Wu et al. (Wu et al., 2024a) constructs a dataset for aligning LLMs with individual preferences by initially creating a diverse pool of 3,310 user personas, which are expanded through iterative self-generation and filtering. This method is similar to PLUM (Magister et al., 2024) that both simulate dynamic interactions through multi-turn conversation trees, allowing LLMs to infer and adapt to user preferences. To enable LLMs to adapt to individual user preferences without re-training, Lee et al. (Lee et al., 2024) utilizes diverse system messages as meta-instructions to guide the models’ behavior. To support this, the MULTIFACETED COLLECTION dataset is created, comprising 197k system messages that represent a wide range of user values. To facilitate real-time, privacy-preserving personalization on edge devices while addressing data privacy, limited storage, and minimal user disruption, Qin et al. (Qin et al., 2024) introduces a self-supervised method that efficiently selects and synthesizes essential user data, improving model adaptation with minimal user interaction.
Research efforts are also increasingly concentrating on developing datasets that assess models’ comprehension of personalized preferences. Kirk et al. (Kirk et al., 2024) introduces PRISM Alignment Dataset that maps the sociodemographics and preferences of 1,500 participants from 75 countries to their feedback in live interactions with 21 LLMs, focusing on subjective and multicultural perspectives on controversial topics. PersonalLLM (Zollo et al., 2024) introduces a novel personalized testdb, which curates open-ended prompts and multiple high-quality responses to simulate diverse latent preferences among users. It generates simulated user bases with varied preferences from pre-trained reward models, addressing the challenge of data sparsity in personalization. ALOE (Wu et al., 2024a) further generates tree-structured multi-turn conversations instead of single-turn pairwise responses.
@@ 5.2 Personalized Alignment Optimization
| Training-Time Personalization | Decoding-Time Personalization | |
| Training Stage | (for each in ) | |
| Inference Stage | (Merging) (Ensemble) | |
| Pros. | Strong Personalization Efficient Inference | High Flexibility No Retraining |
| Cons. | High Training Cost Lack Flexibility | Storage Overhead High Inference Cost |
@@
Personalized preference alignment is usually modeled as a multi-objective reinforcement learning (MORL) problem, where personalized preference is determined as the user-specific combination of multi-preference dimensions. Based on this, a typical alignment paradigm is training-time personalization. This involves using a personalized reward derived from multiple reward models to guide the training of policy LLMs, which is illustrated on the left of Figure 7. MORLHF (Wu et al., 2023) separately trains reward models for each dimension and retrains the policy language models using proximal policy optimization, guided by a linear combination of these multiple reward models. This approach allows for the reuse of the standard RLHF pipeline. MODPO (Zhou et al., 2023) introduces a novel RL-free algorithm extending Direct Preference Optimization (DPO) for managing multiple alignment objectives. It integrates linear scalarization into the reward modeling process, enabling the training of LMs using a margin-based cross-entropy loss as implicit collective reward functions.
@@
The alternative paradigm is decoding-time personalization, which combines multiple trained policy LLMs during inference. Table 4 provides a detailed comparison of this approach against training-time personalization. Its main strategies, personalized weight merging and personalized model ensemble, are illustrated in Figure 7(b) and (c), respectively. Personalized Soups (Jang et al., 2023) and Reward Soups (Rame et al., 2024) address the challenge of RL from personalized human feedback by first training multiple policy models with distinct preferences independently and then merging their parameters post-hoc during inference. Both methods allow for dynamic weighting of the networks based on user preferences, enhancing model alignment and reducing reward misspecification. Also, the personalized fusion of policy LLMs can be achieved not only through parameter merging but also through model ensembling. MOD (Shi et al., 2024) outputs the next token from a linear combination of all base models, allowing for precise control over different objectives by combining their predictions without the need for retraining. The method demonstrates significant effectiveness when compared to the parameter-merging baseline. PAD (Chen et al., 2024a) leverages a personalized reward modeling strategy to generate token-level rewards that guide the decoding process, enabling the dynamic adaptation of the base model’s predictions to individual preferences.
However, these approaches are limited by their reliance on predefined personalization dimensions and a small set of user preferences, which fail to capture the diversity of real-world personalization needs. Recent work has therefore sought to extend reward learning to better accommodate user-specific preferences. For example, VPL (Poddar et al., 2024) employs a variational encoder to encode a small number of preference annotations from a given user into a latent variable that captures individual taste. A reward model is then conditioned on the latent variable and trained by maximizing the evidence lower bound, enabling accurate prediction of rewards aligned with the user’s unique preferences. PREF (Shenfeld et al., 2025) models each user’s personalized reward function as a linear combination of shared base reward functions, which are first learned from multi-user data via matrix factorization. Using an active learning strategy, it can efficiently infer a new user’s combination weights with only a handful of feedback examples (e.g., 10), enabling user-specific adaptation of LLMs’ outputs without retraining. PPT (Lau et al., 2024) leverages in-context learning for scalable personalization by generating two candidate responses for each user prompt, soliciting user rankings, and dynamically incorporating this feedback into the model’s context to adapt to individual preferences over time. Drift (Kim et al., 2025b) proposes a training-free method where user preference is modeled as a linear combination of interpretable attributes, and then computes a reward signal by comparing log-likelihoods under attribute-specific prompts to guide the generation of a frozen base model. In contrast to these offline methods based on static preferences, other approaches focus on real-time adaptation. AMULET (Zhang et al., 2025g) approaches this from another unique test-time perspective, by formulating the decoding process of each token as an independent online learning problem. It obtains the optimization direction by contrasting the model’s output with and without user-provided prompts and utilizes an efficient closed-form solution for real-time iterative optimization, thus adapting to personalized user needs instantly without retraining the model. RLPA (Zhao et al., 2025b) simulates real users via a dynamically updated user portrait model, enabling the language model agent to engage in multi-turn interactive reinforcement learning within this simulated environment, thus continuously adapting its policy to evolving user preferences. PersonaAgent (Zhang et al., 2025a) extends personalization from text style to decision behaviors such as tool usage, guided by a dynamic Persona that directs the agent’s decisions. Using test-time user-preference alignment, it compares the agent’s responses with the user’s real actions in real time, using textual loss as feedback to optimize the Persona, which ensures personalization run through every step of the tool decision.